OpenStudy (anonymous):

Zora saved 9/20 of her paycheck but needed to deduct 25.35 for a bill.If the total that went into her savings was 95.23 what was the amount of zoras paycheck?

OpenStudy (texaschic101):

(9/20)x - 25.35 = 95.23 (9/20)x = 95.23 + 25.35 (9/20)x = 120.58 divide both sides by 9/20 x = 120.58 * 20/9 x = 2411.6/9 x = 267.96 check... (9/20)(267.96) - 25.35 = 95.23 2411.64/20 - 25.35 = 95.23 120.58 - 25.35 = 95.23 95.23 = 95.23 (correct) total paycheck was 267.96